# Making Parametric Anomaly Detection on Tabular Data Non-Parametric Again

### v1

@@ 5.3. Location of the retrieval module
Table 4. Comparison of transformer+attention-bsim across values of $\lambda$
@@ 5.4. Random mask bank
Table 5. Comparison of the performance of the transformer + attention-bsim model for two mask bank set-ups. The same model was used for inference in both set-ups. We report an average over $10$ different splits of the data.
@@ 6. Limitations and Conclusion
As with most non-parametric models, our retrieval-augmented models display a higher complexity than parametric approaches. These approaches can scale well for datasets with a reasonable number of features $d$; however, for large values of $d$, these models involve a high memory cost.
@@ Conclusion
In this work, we have proposed an extensive investigation into external retrieval to augment reconstruction-based anomaly detection methods for tabular data. We have shown that augmenting existing AD methods using attention-based retrieval modules can help foster performance by allowing the model to attend to sample-sample dependencies. Indeed, our experiments involving an extensive benchmark of tabular datasets demonstrate the effectiveness of retrieval-based approaches since the architecture involving the attention-bsim module surpasses the vanilla transformer by a significant margin. Overall, this work emphasizes the pertinence of leveraging sample-sample dependencies to detect anomalies on tabular datasets.
@@
Overall, our work showed that the best-performing attention-based retrieval mechanism relies on other forms of attention than vanilla attention. Parallel to that, models like NPT-AD Thimonier et al. (2023a) have shown strong performance for anomaly detection on tabular data and rely on standard multi-head self-attention through the Attention Between Datapoints (ABD) mechanism, close to the v-attention module, to leverage inter-sample relations. Our findings may invite further research on modifying the ABD mechanism involved in NPTs to improve their AD performance. Finally, the use of external retrieval modules proved effective for the task of anomaly detection using mask reconstruction. The proposed external retrieval module could also be easily added to existing deep-AD methods to test whether they may prove relevant for other pretext tasks for anomaly detection on tabular data.
@@ A.1. Dataset characteristics
A.2. Experimental settings
A.2.1. Training pipeline
  • Masking: We sample from a Bernouilli distribution with probability $p_{mask}$ whether each of the $d$ features is masked.
    $\mbox{mask}=(m_{1},\dots,m_{d}),$
  • Encoding: For numerical features, we normalize to obtain 00 mean and unit variance, while we use one-hot encoding for categorical features. At this point, each feature $j$ for $j\in[1,2,...,d]$ has an $e_{j}$-dimensional representation, $encoded(\mathbf{x}_{j})\in\mathbb{R}^{e_{j}}$ , where $e_{j}=1$ for numerical features and for categorical features $e_{j}$ corresponds to its cardinality. We then mask each feature according to the sampled mask vector and concatenate each feature representation with the corresponding mask indicator function. Hence, each feature $j$ has an $(e_{j}+1)$-dimensional representation
    $((1-m_{j})\cdot encoded(\mathbf{x}_{j}),m_{j})\in\mathbb{R}^{e_{j}+1},$
    where $\mathbf{x}_{j}$ is the $j$-th features of sample $\mathbf{x}$.
  • In-Embedding: We pass each of the features encoded representations of sample $\mathbf{x}$ through learned linear layers $\texttt{Linear}(e_{j}+1,e)$. We also learn $e$-dimensional index and feature-type embeddings as proposed in (Kossen et al., 2021). Both are added to the embedded representation of sample $\mathbf{x}$. The obtained embedded representation is thus of dimension $d\times e$
    $e_{\mathbf{x}}=(e_{\mathbf{x}_{1}},e_{\mathbf{x}_{2}},\dots,e_{\mathbf{x}_{d}})\in\mathbb{R}^{d\times e},$
    and $e_{\mathbf{x}_{1}}\in\mathbb{R}^{e}$.
  • Transformer Encoder: The embedded representation obtained from the previous step is then passed through a transformer encoder. The output of the transformer $e_{\mathbf{x}}$ is of dimension $d\times e$.
  • Out-Embedding: The output of the transformer, $e_{\mathbf{x}}\in\mathbb{R}^{d\times e}$ is then used to compute an estimation of the original $d$-dimensional vector. To obtain the estimated feature $j$, we take the $j$-th $d$-dimensional representation which is output by the transformer encoder, $e_{\mathbf{x}_{j}}\in\mathbb{R}^{d}$, and pass it through a linear layer $\texttt{Linear}(e,e_{j})$, where $e_{j}$ is $1$ for numerical features and the cardinality for categorical features.
  • External Retrieval Modules: During training, for a batch $\mathcal{B}$ composed of $b$ samples, for each sample $\mathbf{x}\in\mathcal{B}$, the entire batch serves as the candidates $\mathcal{C}$. In inference, a random subsample of the training set is used as $\mathcal{C}$. Both for training and inference, when possible memory-wise, we use as $\mathcal{B}$ and $\mathcal{C}$ the entire training set.
    As input, the retrieval module receives a $\mathbb{R}^{d\times e}$ representation for each sample. Operations described in eq. 5, 6, 7 and 8 are performed on the flattened representations of samples $\mathbf{x}_{flatten}\in\mathbb{R}^{d\cdot e}$. After selecting $\mathcal{H}\subseteq\mathcal{C}$, each sample is transformed back to its original dimension to allow aggregation as described in equation 9.
A.2.2. Hyperparameters
For the vanilla transformer, as discussed in section 4, we chose a transformer model comprised of either $2$ or $4$ layers with $4$ attention heads for each model, whether it includes a retrieval module or not. We display in table 7 the varying parameters between datasets for the transformer model. Let us mention that minor changes between retrieval modules may have been included to account for memory requirements (especially for the attention-bsim-bval retrieval module, which we found to be more memory costly). We implemented early stopping with various number of epochs depending on the datasets, we refer the reader to the code made available online for more detail on each run. This lead to a different number of epoch for different seed, we thus do not display in the following table the number of epochs. Overall, learning rate, masking probability, $r$ the maximum number of features masked simultaneously, and $e$ the hidden dimension was kept unchanged across architecture.
Table 7. Datasets hyperparameters for the transformer-based approaches. When the batch size is $-1$ it refers to a full pass over the training set before an update of the weights. Due to memory requirements, for some retrieval-based approaches, the batch size may have been adjusted. The column card$(\mathcal{C})$ corresponds to the cardinality of the random subsamples of the training set used to select the $\mathcal{H}$ during inference. During training, $\mathcal{C}$ corresponds to the batch and has the batch size as cardinality. We refer the reader to the code made available online for more details on the hyperparameters.
Dataset num layers batch size lr $p^{train}_{mask}$ $r$ $m$ $e$ card$(\mathcal{C})$
Wine $2$ $-1$ $0.001$ $0.15$ $3$ $78$ $8$ $-1$
Lympho $2$ $-1$ $0.001$ $0.15$ $1$ $18$ $16$ $-1$
Glass $2$ $-1$ $0.001$ $0.15$ $1$ $9$ $16$ $-1$
Vertebral $2$ $-1$ $0.001$ $0.15$ $1$ $6$ $8$ $-1$
WBC $2$ $-1$ $0.01$ $0.15$ $1$ $30$ $16$ $-1$
Ecoli $2$ $-1$ $0.001$ $0.15$ $1$ $7$ $16$ $-1$
Ionosphere $2$ $-1$ $0.001$ $0.15$ $1$ $33$ $8$ $-1$
Arrhythmia $2$ $-1$ $0.001$ $0.15$ $1$ $274$ $8$ $-1$
BreastW $4$ $-1$ $0.001$ $0.15$ $1$ $9$ $8$ $-1$
Pima $4$ $-1$ $0.001$ $0.15$ $4$ $162$ $8$ $-1$
Vowels $4$ $-1$ $0.001$ $0.15$ $2$ $78$ $16$ $-1$
Letter Recognition $4$ $-1$ $0.001$ $0.25$ $1$ $32$ $16$ $-1$
Cardio $4$ $-1$ $0.001$ $0.15$ $2$ $231$ $16$ $-1$
Seismic $4$ $-1$ $0.001$ $0.25$ $1$ $11$ $16$ $-1$
Musk $4$ $-1$ $0.001$ $0.15$ $1$ $166$ $8$ $-1$
Speech $4$ $128$ $0.001$ $0.15$ $1$ $400$ $8$ $64$
Thyroid $4$ $-1$ $0.001$ $0.15$ $3$ $41$ $8$ $-1$
Abalone $4$ $-1$ $0.001$ $0.15$ $2$ $45$ $16$ $-1$
Optdigits $4$ $2048$ $0.001$ $0.15$ $1$ $64$ $16$ $1024$
Satimage-2 $4$ $2048$ $0.001$ $0.2$ $1$ $36$ $16$ $1024$
Satellite $4$ $2048$ $0.001$ $0.15$ $1$ $36$ $16$ $1024$
Pendigits $4$ $-1$ $0.001$ $0.15$ $1$ $16$ $16$ $-1$
Annthyroid $4$ $-1$ $0.001$ $0.15$ $1$ $6$ $16$ $-1$
Mnist $4$ $-1$ $0.001$ $0.15$ $1$ $100$ $8$ $256$
Mammography $4$ $-1$ $0.001$ $0.15$ $1$ $6$ $16$ $1024$
Shuttle $4$ $2048$ $0.001$ $0.15$ $3$ $129$ $16$ $1024$
Mulcross $4$ $2048$ $0.001$ $0.15$ $1$ $4$ $16$ $1024$
ForestCover $4$ $1024$ $0.001$ $0.15$ $2$ $55$ $16$ $1024$
Campaign $4$ $2048$ $0.001$ $0.15$ $1$ $62$ $16$ $512$
Fraud $4$ $4096$ $0.001$ $0.2$ $1$ $29$ $32$ $1024$
Backdoor $4$ $256$ $0.001$ $0.2$ $1$ $196$ $32$ $512$
@@ B.1. Detailed tables for main experiments
B.2. Comparison to existing models
(a) F1-score ($\uparrow$)
(b) Rank ($\downarrow$)
Figure 2. For each of the $31$ datasets on which models were evaluated,we report the average F1-score over $20$ runs for $20$ different seeds. See (Thimonier et al., 2023a) for detail on the obtained metrics. For both figures, the model displayed on the far left is the worst-performing model for the chosen metric, and on the far right is the best-performing model. We also highlight the metric of the best-performing model in bold.
For completeness we compare the different architectures proposed in the present work to existing methods in the literature. We rely on the experiments conducted in (Thimonier et al., 2023a) and Shenkar and Wolf (2022) for the metrics of the competing methods. We display in figure 2 the comparison to existing methods. We compare our methods to recent deep methods, namely GOAD (Bergman and Hoshen, 2020), DROCC (Goyal et al., 2020), NeuTraL-AD (Qiu et al., 2021), the contrastive approach proposed in (Shenkar and Wolf, 2022) and NPT-AD (Thimonier et al., 2023a). We also compare to classical non-deep methods such as Isolation Forest (Liu et al., 2008), KNN (Ramaswamy et al., 2000), RRCF (Guha et al., 2016), COPOD (Li et al., 2020) and PIDForest (Gopalan et al., 2019). We refer the reader to Thimonier et al. (2023a) for the detail on the F1-score per dataset for other methods than those shown in tables 8 and 9.

### v2

@@ 5.4. Location of the retrieval module
Table 5. Comparison of transformer+attention-bsim across values of $\lambda$ based on the F1-Score ($\uparrow$).
@@ 5.5. Random mask bank
Table 6. Comparison of the performance of the transformer + attention-bsim model based on the F1-Score ($\uparrow$) for two mask bank set-ups. The same model was used for inference in both set-ups. We report an average over $10$ different splits of the data.
@@ 6. Limitations and Conclusion
As with most non-parametric models that leverage the training set in inference, our retrieval-augmented models display a higher complexity than parametric approaches. These approaches can scale well for datasets with a reasonable number of features $d$; however, for large values of $d$, these models incur a high memory cost.
@@ Conclusion
In this work, we have proposed an extensive investigation into external retrieval to augment reconstruction-based anomaly detection methods for tabular data. We have shown that augmenting existing AD methods using attention-based retrieval modules can help foster performance by allowing the model to attend to sample-sample dependencies. Indeed, our experiments involving an extensive benchmark of tabular datasets demonstrate the effectiveness of retrieval-based approaches since the architecture involving the attention-bsim module surpasses the vanilla transformer by a significant margin. We also provide a first explanation as to why combining both types of dependencies can be critical to obtaining consistent performance across datasets: different types of anomalies require different types of dependencies to be efficiently detected.
@@
Overall, our work showed that the best-performing attention-based retrieval mechanism relies on other forms of attention than vanilla attention. Parallel to that, models like NPT-AD have shown strong performance for anomaly detection on tabular data and rely on standard multi-head self-attention through the Attention Between Datapoints (ABD) mechanism, close to the v-attention module, to leverage inter-sample relations. Our findings may invite further research on modifying the ABD mechanism involved in NPTs to improve their AD performance. Finally, the use of external retrieval modules proved effective for the task of anomaly detection using mask reconstruction. The proposed external retrieval module could also be easily added to existing deep-AD methods to test whether they may prove relevant for other pretext tasks for anomaly detection on tabular data.
Acknowledgment
This work was granted access to the HPC resources of IDRIS under the allocation 2023-101424 made by GENCI. This research publication is supported by the Chair ”Artificial intelligence applied to credit card fraud detection and automated trading” led by CentraleSupelec and sponsored by the LUSIS company.
@@ B.1. Detailed tables for main experiments
B.2. Synthetic dataset generation and experimental detail for section 5.1
The synthetic three dimensional dataset was generated as follows.
  • Normal samples: We consider an interval of values $[-2,3]$ from which we uniformly sample the first feature $x_{1}$. We then sample $x_{2},x_{3}$ following the relation given in eq. (12) with parameters, $(\alpha_{1},\beta_{1})=(2,3)$, $(\alpha_{2},\beta_{2})=(4,3)$ and $\varepsilon\sim\mathcal{N}(0,1)$.
  • Anomalies (type 1): We consider an interval of values $[3.3,4]$ from which we uniformly sample the first feature $x_{1}$ and keep the rest as for normal samples.
  • Anomalies (type 2): We consider an interval of values $[1.5,2.5]$ from which we uniformly sample the first feature $x_{1}$ and sample $x_{2},x_{3}$ following eq. (12) but with parameters $(\alpha_{1},\beta_{1})=(-7.5,-1)$ and $(\alpha_{2},\beta_{2})=(4,3)$.
The vanilla transformer and its augmented version were trained with the following hyperparameters:
  • Batch size: $-1$.
  • Patience: $100$ epochs.
  • Learning rate (lr): $0.001$.
  • Hidden dim ($e$): $8$.
  • Masking probability $p_{mask}$: $0.15$.
  • Number of attention heads: $2$.
  • Number of layers of the encoder: $2$.
  • Retrieval hyper-parameters:
    • Retrieval location: post-encoder
    • Retrieval aggregation location: post-encoder
    • $\lambda$: 0.5
    • $\mathcal{C}=\mathcal{D}_{train}$
    • $card(\mathcal{H})=500$
For Mask-KNN, following (Thimonier et al., 2024) we set the number of neighbors to $k=5$.